OpenStudy (anonymous):

i need help pleez! how do you solve sin^2x + cosx=2

OpenStudy (anonymous):

i find that it is a no solution problem but i dont think that can be right..

OpenStudy (anonymous):

recall that \[\sin ^{2} x = 1- \cos ^{2}\] so now we have \[1 -\cos ^{2}x +cosx = 2\] Bring the 2 over by subtracting from both sides let y = cosx \[-y ^{2} +y +1\] can you factor that...

OpenStudy (anonymous):

Woops I meant\[-y ^{2} +y -1\] haha.. yea factor this.

OpenStudy (anonymous):

it isnt factorable right?

OpenStudy (anonymous):

it is...try